alexi throws a pebble into a unused well. the distance D,the pebble falls after t seconds is given by the equation d=16tsquared + 56T, where d is measured in feet. if the water in the well is 240 feet below ground level,how many secconds will it take for the pebble to hit the water?
Answer by stanbon(75887)

alexi throws a pebble into a unused well. the distance D,the pebble falls after t seconds is given by the equation d=16tsquared + 56T, where d is measured in feet. if the water in the well is 240 feet below ground level,how many secconds will it take for the pebble to hit the water?
--------------------------------------------------------
16t^2+56t = 240
2t^2 + 7t = 30
t^2 + 7t - 30 = 0
(t+10)(t-3) = 0
Positive solution:
t = 3 seconds
==========================
Cheers,
Stan H.